Mike and Jake are brothers. They are both younger than 10. Mike is 4 times older than Jake. In 4 years, Mike will be twice as old as Jake. How old are Mike and Jake?
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

Mike's age = M
Jake's age = J

M = 4J
M + 4 = 2(J + 4)

Plug in the first equation into the second.

4J + 4 = 2J + 8
2J = 4

Jake is 2 and Mike is 8.

Answer by ikleyn(52790)

.

You need to read from this source to learn how to use properly

"three times older than . . . " or "three times as old as . . . "


https://forum.wordreference.com/threads/times-older-than-times-as-old-as.3556657/